OpenStudy (4meisu):

Consider the equation 3 cos 2x + sin x = 1 (a) Write this equation in the form f(x) = 0, where f(x) = a sin^2 x + b sin x + c (b) factor f(x) (c) Write down the number of solutions of f(x) = 0, for 0 < x <2π (do not solve 0)

OpenStudy (anonymous):

Use the formula \[\cos(2 x) = \cos^2(x) - \sin^2(x) = 1- 2 \sin^2(x) \]

OpenStudy (anonymous):

\[3 \cos (2x) + \sin( x)= 1\\ 3( 1- 2 \sin^2(x)) + \sin(x) -1=0\\ 3 - 6 \sin^2(x) + \sin(x) -1 =0\\ -6 \sin^2(x) + \sin(x) + 2 =0\\ 6 \sin^2(x) - \sin(x) - 2 =0\\ (2 \sin (x)+1) (3 \sin (x)-2)= \]

OpenStudy (anonymous):

You should be able to finish it now.